### v1

# HEIST: A Graph Foundation Model for Spatial Transcriptomics and Proteomics Data

Abstract
Single-cell transcriptomics has become a great source for data-driven insights into biology, enabling the use of advanced deep learning methods to understand cellular heterogeneity and transcriptional regulation at the single-cell level. With the advent of spatial transcriptomics data we have the promise of learning about cells within a tissue context as it provides both spatial coordinates and transcriptomic readouts. However, existing models either ignore spatial resolution or the gene regulatory information. Gene regulation in cells can change depending on microenvironmental cues from neighboring cells, but existing models neglect gene regulatory patterns with hierarchical dependencies across levels of abstraction. In order to create contextualized representations of cells and genes from spatial transcriptomics data, we introduce HEIST, a hierarchical graph transformer-based foundation model for spatial transcriptomics and proteomics data. HEIST models tissue as spatial cellular neighborhood graphs, and each cell is, in turn, modeled as a gene regulatory network graph. The framework includes a hierarchical graph transformer that performs cross-level message passing and message passing within levels. HEIST is pre-trained on 22.3M cells from 124 tissues across 15 organs using spatially-aware contrastive learning and masked auto-encoding objectives. Unsupervised analysis of HEIST representations of cells, shows that it effectively encodes the microenvironmental influences in cell embeddings, enabling the discovery of spatially-informed subpopulations that prior models fail to differentiate. Further, HEIST achieves state-of-the-art results on four downstream task such as clinical outcome prediction, cell type annotation, gene imputation, and spatially-informed cell clustering across multiple technologies, highlighting the importance of hierarchical modeling and GRN-based representations.
1 Introduction
Single-cell RNA sequencing (scRNA-seq) technologies have revolutionized our ability to study gene expression at the resolution of individual cells, offering data-driven insights into biology. The complexity of single-cell data has fueled the development of machine learning [7, 20, 16, 39, 3, 8] methods for modeling cellular diversity, uncovering regulatory mechanisms, predicting cell states, and imputing genes. However, scRNA-seq measurements are unaware of the spatial context of cells within tissues, which is critical for understanding processes such as tissue organization, microenvironment interactions, and how gene regulation gives rise to tissue-level behaviors. Spatial transcriptomics is an emerging technology that bridges this gap by preserving the physical locations of gene expression measurements, enabling comprehensive studies of tissue architecture, cell-cell communication, and tumor microenvironments [42, 31]. Although spatial transcriptomics technology has advanced rapidly, current datasets are small due to low-throughput of sequencing technologies, vary across platforms, organs, and technologies, and often lack detailed labels for supervised learning. These issues requires specialized models to be trained for every new dataset or organ challenging. A foundation model trained on diverse spatial transcriptomics data overcomes these challenges by learning a generalized representation across different tissues, organs, experimental settings, and technologies. This enables application downstream tasks to be performed more effectively, even when data is limited.
While recent foundation models such as scGPT [7] and scFoundation [18] have extended large language models to scRNA-seq data, they do not explicitly model cell-cell relationships. Capturing these cell-cell interactions is vital, as the cellular microenvironment plays a key role in regulating cell states, immune infiltration, tissue development, and tumor progression [22, 6]. In response, graph-based methods such as GraphST [26] and STAGATE [9] have leveraged spatial graphs to model local cell neighborhoods. However, these methods are primarily task-specific and do not support general-purpose, transferable representation learning, limiting their ability to adapt across datasets and organs. CellPLM [39] and scGPT-spatial [7] are foundation models that incorporates cell-cell interactions. But they are pre-trained on a pre-defined set of genes, hindering their generalization to unseen genes. Crucially, none of these approaches capture the influence of gene regulatory networks (GRNs), which drive cell function [28]. GRNs do not act in isolation—they are shaped by, and in turn shape, the spatial microenvironment [49]. These hierarchical relationships between GRNs and cells are fundamental to biological organization [21, 25, 37], ignoring which means missing how gene programs in one cell influence neighboring cells and contribute to tissue-level phenomena [32].
To this end, we propose HEIST (Hierarchical EmbeddIngs for Spatial Transcriptomics), the first spatial transcriptomics foundation model that explicitly models both cell-cell spatial interactions and GRNs, while enabling cross-modal information exchange between cells and genes. In HEIST, cell is associated with a distinct GRN, constructed from co-expression relationships, enabling cross-cell influence between GRNs: as cells interact through spatial graphs, their local microenvironments modulate gene regulation. By explicitly modelling GRNs, HEIST connects gene-level interactions with tissue-level phenotypes, unlocking deeper biological insights that prior models miss. Prior hierarchical graph neural networks [14, 12, 15, 4] either lack mechanisms for cross-level information exchange or only operate on hierarchical graphs within a single data modality (e.g., different levels of resolution in a mesh). HEIST addresses both key challenges of hierarchical modeling: it performs cross-level message passing between genes and cells to integrate spatial and regulatory information. Specifically, gene embeddings are updated using information from their parent cell’s spatial context, while cell embeddings are refined by aggregating their constituent gene representations. As a result, the same gene can acquire different embeddings depending on its spatial context, capturing how microenvironmental factors influence gene regulation, enabling richer representations.
Figure 1: Overview of the HEIST framework. (A) HEIST is pre-trained on a large-scale spatial transcriptomics dataset spanning 124 tissues and 15 organs (22.3M cells). (B) HEIST encodes both gene regulatory networks (GRNs) and spatial cell graphs to support downstream tasks such as cell clustering, gene imputation, and clinical outcome prediction (e.g., immunotherapy response, remission). The HEIST decoder can be fine-tuned while the encoder remains frozen.
We pretrain HEIST (Figure 1(A)) on a large-scale corpus of spatial transcriptomics data comprising over 22.3 million cells from 124 tissues across 15 organs and two technologies. HEIST is evaluated on four downstream tasks (Figure 1(B))—clinical outcome prediction, cell type annotation, gene imputation, and cell clustering—achieving state-of-the-art performance across seven organs. In addition to improved predictive accuracy, HEIST enables the discovery of spatially-informed cellular subpopulations by capturing microenvironmental influences that prior models fail to differentiate. HEIST also offers significant computational efficiency compared to models that learn gene representations, being 8× faster than scGPT-spatial and 48× faster than scFoundation. These results highlight HEIST’s ability to generalize across tissues, technologies, and even unseen genes, without requiring extensive fine-tuning.
We summarize out contributions as follows:
  • Modeling inter-cellular and tissue-scale effects of GRNs: HEIST is the first foundation model for spatial transcriptomics to explicitly incorporate GRNs alongside spatial graphs, enabling a local gene programs to influence tissue-level organization, and vice versa.
  • Hierarchical representation learning with biological inductive bias: Using biologically motivated hierarchical modeling, HEIST captures fine-grained gene regulation within cells and long-range cellular interactions through novel cross-level message passing mechanism, producing adaptive, biologically contextualized embeddings.
  • A task-agnostic, general-purpose foundation model: HEIST is trained in a self-supervised manner on a large-scale corpus of spatial transcriptomics data comprising over 22.3 million cells spanning 15 organs and 124 tissues types. In downstream evaluations, HEIST achieves state-of-the-art performance across four diverse tasks—outperforming prior models, while being computationally efficient.
2 Related works
This section outlines hierarchical graph neural networks and their limitations, followed by graph-based spatial transcriptomics methods and foundation models for single-cell and spatial data.
Hierarchical graph neural networks. Hierarchical graph neural networks operate on graphs with multiple levels, where each level represents a distinct granularity or type of information. MSMGN [12], HOOD [15], and BSMS-GNN [4] perform cross-level message passing across levels, but the levels are of same modality at different resolutions (e.g., coarsened versions of spatial graphs), rather than integrating graphs of distinct modalities and they are geared towards propagating signals globally. HIGH-PPI [14] introduces a hierarchical graph where one level encodes protein-protein interactions and the other encodes residue-level interactions. However, it does not include cross-level message passing, which limits the ability to combine information across biological levels. In contrast, HEIST works information from different modalities while integrating information across levels.
Graph-based Spatial Transcriptomics methods. Graph-based spatial transcriptomics methods construct cell graphs based on spatial proximity and apply graph neural networks to model spatial relationships. SpaGCN [20], CCST [24], STAGATE [9], ConST [48], and GraphST [26] train graph neural networks in self-supervised or unsupervised manner by using techniques such as deep graph infomax (DGI) [36], iterative clustering [2], and auto-encodings task on gene expression [11]. Unlike HEIST, these models focus on specific downstream tasks such as cell clustering and do not generalize to other tissues without retraining. Furthermore, they do not incorporate hierarchical modeling across genes and cells.
scRNA-seq and Spatial Transcriptomics Foundation models. Foundation models for single-cell and spatial transcriptomics aim to pretrain general-purpose representations that can be transferred across datasets and tasks. scFoundation [18], geneformer [33], and scGPT [7] extend transformer architectures to model scRNA-seq data by treating gene expression profiles as sequences. However, these models assume an ordering over genes and do not explicitly capture cell-cell relationships or spatial information. scGPT-spatial [38] adapts scGPT to spatial transcriptomics. CellPLM [39] incorporates cell-cell interactions in pertaining and adds a gaussian mixture prior to overcome data limitations. But CellPLM and scGPT-spatial operate over a fixed set of genes and do not model hierarchy. Existing foundation models thus either neglect spatial dependencies, gene regulatory structure, or both, limiting their ability to capture the context-specific nature of spatial transcriptomics. In contrast, HEIST hierarchically models both this information and computes expressive and generalizable representations by working over GRNs.
3 Method
In this section, we describe the architecture of HEIST, a hierarchical graph transformer designed to learn multi-level embeddings for spatial transcriptomics data. HEIST takes as input a set of graphs $\{{\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\}$, where ${\mathcal{G}}_{c}$ is a spatial graph capturing cell-cell interactions, ${\mathcal{C}}$ is the set of cells, ${\mathcal{E}}$ is the set of spatial edges, ${\mathbf{P}}\in\mathbb{R}^{|{\mathcal{C}}|\times 2}$ represents the spatial positions, and ${\mathcal{T}}$ is the set of cell types. Each graph ${\mathcal{G}}_{g}^{k}$ represents a GRN within cell $k$, capturing gene-gene interactions specific to that cell; ${\mathcal{V}}$ is the set of genes, and ${\mathcal{E}}_{k}$ and ${\mathbf{X}}_{k}$ denote the edges of GRN and gene expression values for cell $k$, respectively. We explain the graph-creation process in Appendix Section A. Given these graphs, HEIST computes cell and gene embeddings ${\mathbf{Z}}_{c}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ and ${\mathbf{Z}}_{g}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$, such that
${\mathbf{Z}}_{c},{\mathbf{Z}}_{g}=\text{HEIST}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\right).$
The model first computes positional embeddings. As shown in Figure 2, model performs intra-level message passing within each graph, followed by cross-level message passing to integrate multi-modal information. HEIST is pre-trained using a combination of contrastive and auto-encoding objectives on gene expression and cell locations. By using these components, HEIST can learn expressive and context-aware cell and gene embeddings that reflect biologically meaningful relationships between cells and genes.
Figure 2: HEIST Architecture
3.1 HEIST Architecture.
The input cell embeddings ${\mathbf{H}}_{c}^{(0)}$ and gene embeddings for cell $k$ ${\mathbf{H}}_{g}^{k^{(0)}}$ are passed through HEISTLayer $L$ times, and the representations are calculated using the equation below:
${\mathbf{H}}_{c}^{(l)},{\mathbf{H}}_{g}^{(l)}=\texttt{HEISTLayer}({\mathbf{H}}_{c}^{(l-1)},{\mathbf{H}}_{g}^{(l-1)},{\mathcal{E}},\{{\mathcal{E}}_{k}\}_{k=0}^{|{\mathcal{C}}|})$
HEISTLayer is divided into two steps, intra-level message passing and cross-level message passing. We perform intra-message passing and calculate the intermediate representations as shown in equation below:
$\tilde{{\mathbf{H}}}_{c}^{(l)}=\texttt{CellGraphTransformer}({\mathbf{H}}_{c}^{(l-1)},{\mathcal{E}}),\tilde{{\mathbf{H}}}_{g}^{(l)}=\texttt{GeneGraphTransformer}({\mathbf{H}}_{g}^{(l-1)},\{{\mathcal{E}}_{k}\}_{k=0}^{|{\mathcal{C}}|})$
where CellGraphTransformer and GeneGraphTransformer are graph transformers as explained in the Appendix Section B.2.
Cross-level message passing. To integrate spatial and gene modalities, HEIST performs cross-level message passing between cell and gene graphs at each layer (Figure 2). Gene embeddings are updated based on spatial context through their parent cell’s embedding, while cell embeddings are refined using pooled summaries of their genes, ensuring transcriptional states shape spatial identity. This bidirectional interaction captures tissue hierarchy, where gene expression depends on both regulatory signals and spatial microenvironments. HEIST thus learns representations that reflect both local gene regulation and large-scale tissue structures.
We use a directional attention mechanism to perform cross-level message passing between gene and cell graphs. Let $\tilde{{\mathbf{H}}}_{g}^{(l)}$ and $\tilde{{\mathbf{H}}}_{c}^{(l)}$ denote the intermediate gene and cell embeddings at layer $l$ after intra-level updates and before cross-level integration. HEIST updates these representations using the equation below:
${\mathbf{H}}_{g}^{(l)}=\texttt{CrossMessagePassingLayer}(\tilde{{\mathbf{H}}}_{g}^{(l)},\tilde{{\mathbf{H}}}_{c}^{(l)^{\text{repeat}}}),\quad{\mathbf{H}}_{c}^{(l)}=\texttt{CrossMessagePassingLayer}(\tilde{{\mathbf{H}}}_{c}^{(l)},\bar{{\mathbf{H}}}_{g}^{(l)}),$
where $\tilde{{\mathbf{H}}}_{c}^{(l)^{\text{repeat}}}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$ represents the cell embeddings repeated $|{\mathcal{V}}|$ times to align with the gene embeddings in each cell. Each gene receives information from its parent cell, enabling spatial context to modulate gene-level representations. Conversely, $\bar{{\mathbf{H}}}_{g}^{(l)}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ is obtained by aggregating the gene embeddings within each cell: $\bar{{\mathbf{H}}}_{g}^{(l)}=[\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{1}),\ldots,\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{|{\mathcal{C}}|})]$, where $\texttt{AGG}(\cdot)$ can be aggregation function such as MEAN pooling or differential pooling [45]. This allows the cell embedding to be informed by the internal transcriptional state of the cell. We use the directional attention mechanism to perform cross message passing as shown in equation below:
[…]
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ are learned weight matrices, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > is the inner product computed row-wise.
Advantages. This formulation enables targeted, direction-aware communication between modalities while preserving their structure and semantics. It is well suited for spatial transcriptomics, where cell and gene representations must be coupled yet retain distinct meanings. Gene expression depends not only on intrinsic regulation but also on spatial context, such as cell interactions and microenvironments. By letting each gene attend to its parent cell embedding, HEIST shapes gene representations in a cell-specific, spatially informed way, and vice versa. Unlike symmetric attention or feature concatenation, this directional mechanism preserves data hierarchy, respects modality roles, and captures how local gene regulation drives tissue organization. Because directional message passing respects the natural hierarchy between genes and cells, allowing each to influence the other without collapsing their distinct biological roles, enabling HEIST to learn spatially informed, biologically grounded representations that generalize across tissues and experiments.
Finally, the intra- and cross-level message passing steps are repeated $L$ times, yielding final embeddings
${\mathbf{Z}}_{c},{\mathbf{Z}}_{g}=\text{HEIST}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\right).$
Decoder. After calculating the final representations, we pass the embeddings into a decoder to reconstruct the original spatial locations using HEIST-Decoder:
$\hat{{\mathbf{P}}},\{\hat{{\mathbf{X}}}_{k}\}_{k=0}^{|{\mathcal{C}}|}=\text{HEIST-Decoder}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k})\}_{k=0}^{|{\mathcal{C}}|},{\mathbf{Z}}_{c},{\mathbf{Z}}_{g}\right).$
where, HEIST-Decoder is a 3-layer GIN network [43].
3.2 Pre-training tasks.
HEIST is trained on a combination of contrastive and masked auto-encoding objectives.
Contrastive. We use a contrastive objective to learn context-aware representations by bringing similar cells and genes—such as neighboring cells of the same type or co-expressed genes—closer in embedding space while pushing dissimilar pairs apart. This separates functionally distinct cell populations and gene modules, even when spatially close. Additionally, we introduce cross-level contrastive alignment to ensure consistency between gene and cell representations, so cell embeddings reflect gene expression patterns and gene embeddings incorporate spatial context. The contrastive loss is calculated using the equation below:
$\ell_{c\leftrightarrow c}=\log\frac{e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})/\tau}}{\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,k})/\tau}},\quad\ell_{g\leftrightarrow g}=\log\frac{e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q})/\tau}}{\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{k}_{g,r})/\tau}},\quad\ell_{c\leftrightarrow g}=\log\frac{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}}{\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{k}_{g})/\tau}}$
$\mathcal{L}_{\text{contrastive}}=(\sum_{i,j\in\mathcal{P}}\ell_{c\leftrightarrow c}({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})+\ell_{c\leftrightarrow g}({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})+\sum_{(p,q)\in{\mathcal{V}}^{2}}\ell_{g\leftrightarrow g}({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q}))$
where $\theta(\cdot,\cdot)$ denotes a similarity function (e.g., cosine similarity), $c\leftrightarrow c,g\leftrightarrow g$, and $c\leftrightarrow g$ denotes contrastive loss between cells, genes, and cell-genes respectively, $\tau$ is a temperature parameter that controls the sharpness of the contrastive distribution, and $t_{i}$ is the cell type label of cell $i$. The set of positive pairs $\mathcal{P}=\{(i,j)\mid t_{i}=t_{j},\,d(i,j)\leq r\}$, where $d(i,j)$ is the spatial distance between cells $i$ and $j$, and $r$ is a fixed spatial radius. Similarly, the set of negative samples for cell $i$ $\mathcal{N}_{i}=\{k\mid t_{i}\neq t_{k},\,d(i,k)\leq r\}$, i.e., cells within radius $r$ that belong to a different cell type.
Masked-auto encoding. We also train HEIST with a masked auto-encoding loss to improve reconstruction and robustness. By masking subsets of cell and gene nodes, the model learns to reconstruct gene expression and spatial coordinates from the remaining context, reflecting real-world challenges like dropout and noise in spatial transcriptomics. This encourages the model to infer missing data, generalize across datasets, and use gene signals to recover spatial context and spatial cues to predict gene expression. After reconstructing the spatial locations and gene-expression, the masked auto-encoding loss is calculated using equation below:
$\mathcal{L}_{\text{mae}}=\text{MSE}(\hat{{\mathbf{P}}}\odot mask_{c},{\mathbf{P}}\odot mask_{c})+\frac{1}{|{\mathcal{C}}|}\sum_{k=0}^{|{\mathcal{C}}|}\text{MSE}(\hat{{\mathbf{X}}}_{k}\odot mask^{k}_{g},{\mathbf{X}}_{k}\odot mask^{k}_{g}),$
where $mask_{c}$ is the mask over the cell locations, and $mask_{g}^{k}$ is the gene-expression mask for cell $k$.
Final loss. Contrastive learning structures the latent space to emphasize biological similarities and differences, promoting better separation of cell types and gene programs. In contrast, masked autoencoding ensures that embeddings retain rich information content necessary for reconstructing gene expression and spatial locations. Together, they prevent trivial or collapsed representations and produce embeddings that are both discriminative and information-rich. Hence, the final objective is a weighted sum of the contrastive and autoencoding losses, along with an orthogonality regularization which encourages the embedding dimensions to be decorrelated, promoting diverse and non-redundant representations [46]:
$\mathcal{L}=\sigma(\gamma)\cdot\mathcal{L}_{\text{contrastive}}+(1-\sigma(\gamma))\cdot\mathcal{L}_{\text{mae}}+\lambda\left(\left\|\mathbf{I}_{d}-\mathbf{Z}_{c}^{\top}\mathbf{Z}_{c}\right\|_{F}^{2}+\left\|\mathbf{I}_{d}-\mathbf{Z}_{g}^{\top}\mathbf{Z}_{g}\right\|_{F}^{2}\right),$
where $\lambda$ is a regularization weight, and $\gamma$ is a learnable scalar that dynamically balances the contrastive and reconstruction terms.
Positional encoding. HEIST incorporates positional encodings (PE) at both the cell and gene levels to inject spatial and regulatory structure into the learned representations. As opposed to traditional graph PEs such has Laplacian PE [10], random walk PE [10], node-centrality based PE [44], we make use of sinusoidal PEs. This design choice is motivated by the fact that sinusoidal PE yields expressive representations and is computationally inexpensive as opposed to traditional graph PEs, leading to efficient and expressive representations [40]. Specifically, we use 2D-sinusoidal embeddings and rank-based sinusoidal described in the Appendix Section B.1. We then add the spatial information and gene-expression to the positional encodings using the equations below:
$\text{PE}_{c}=\texttt{2DSinusoidal}({\mathbf{P}}),\text{PE}^{k}_{g}=\texttt{RankSinusoidal}({\mathbf{X}}_{k});\quad{\mathbf{H}}_{c}^{(0)}=\text{PE}_{c}+{\mathbf{P}},{\mathbf{H}}_{g}^{k^{(0)}}=\text{PE}^{k}_{g}+{\mathbf{X}}_{k}$
where ${\mathbf{H}}_{c}^{(0)}$ and ${\mathbf{H}}_{g}^{k^{(0)}}$ are the input cell and gene embeddings for cell $k$, respectively.
Computational efficiency As shown in Table A4 in the Appendix, HEIST demonstrates significant computational advantages, achieving 8× faster embedding extraction time compared to scGPT-spatial and 48× faster than scFoundation. This efficiency comes from HEIST’s hierarchical modeling and sparse cross-level message passing, which avoids the expensive full self-attention computations required by large language model-based approaches.
4 Experiments
In this section, we first describe the pretraining datasets used to train HEIST with contrastive and autoencoding objectives. We then outline the downstream tasks and corresponding datasets, followed by baselines, results, insights, and ablation studies.
Pretraining Datasets. HEIST is trained on a large and diverse collection of high-resolution spatial transcriptomics datasets, primarily generated using single-cell technologies such as MERFISH and Xenium. The pretraining dataset comprises 22.3 million cells from 124 tissue slices across 15 organs [cf. Figure 1(A)], including 13.3M cells from 10x Genomics  [^1] https://www.10xgenomics.com/, 8.7M from Vizgen  [^2] https://vizgen.com/, and 360k cells from the Seattle Alzheimer’s Brain Atlas  [^3] https://portal.brain-map.org/explore/seattle-alzheimers-disease. The large scale and diversity of the dataset improve the reliability and robustness of learned representations, enabling better transferability to downstream tasks across varying biological contexts and experimental platforms. While integrating data from different technologies presents challenges such as batch effects and varying gene panels, HEIST’s architecture and training objectives are specifically designed to overcome these issues and learn biologically meaningful representations. A detailed breakdown of datasets, descriptions, and sources is provided in Appendix C.
Experimental setup. We detail pretraining hyperparameters in Appendix Table A1. HEIST was pretrained on 4 NVIDIA L40s GPUs (40GB each), with each epoch taking approximately 3 hours. Although the maximum number of epochs was set to 20, early stopping based on validation loss typically halted training around Epoch 5 or 6. For downstream evaluations, we assess HEIST in both zero-shot and fine-tuning settings. In the zero-shot setting, the pretrained model is directly evaluated on unseen spatial transcriptomics datasets to assess generalization without further training. For fine-tuning, we first extract embeddings from the frozen model and either train an MLP prediction head or fine-tune the decoder. We perform each experiment 5 times provide mean and standard deviation, except in Charville and UPMC dataset where we use the split the data using method from Wu et al. [41]. The code is available at https://anonymous.4open.science/r/HEIST-908D.
Downstream Tasks. We evaluate HEIST across four spatial transcriptomics technologies, six organs, and four downstream tasks—cell clustering, cell type annotation, clinical outcome prediction, and gene imputation—to assess both biological insight discovery and clinical relevance in zero-shot and fine-tuned settings.
Cell clustering is critical for discovering novel cell types and understanding how microenvironmental factors shape cellular behavior, particularly in tumor microenvironments. An expressive model should not only cluster cells but also reveal microenvironment-driven subclusters, providing insights into spatially informed cell populations. Clustering is performed using frozen embeddings, and performance is evaluated on datasets such as SEA-AD [13], Charville, UPMC, DFCI [41], and Merfish Lung Cancer [5] using normalized mutual information (NMI). Cell type annotation classifies cells into known biological categories, enabling functional interpretation of cellular diversity. HEIST embeddings are extracted from the frozen model, and an MLP classifier is trained on labeled datasets including SEA-AD, Charville, UPMC, DFCI, and MERFISH lung cancer, with performance evaluated using F1 score.
Clinical outcome prediction aims to classify entire tissues, predicting outcomes such as immunotherapy response, treatment outcomes, remission status, and placenta condition. This task is essential for clinical decision-making and understanding disease progression. HEIST is evaluated on datasets including Charville (Colon), UPMC (Neck), and DFCI (Neck) [41] collected using CODEX, skin cancer data [30] collected using MIBI, and placenta data collected using Xenium. Predictions are made using frozen cell embeddings and an MLP classifier, with AUC-ROC reported as the evaluation metric. Gene imputation recovers missing or noisy gene expression values, a common issue in spatial transcriptomics due to measurement limitations. We perform gene imputation by predicting masked gene values, using stratified sampling based on gene sparsity following the approach of Avşar and Pir [1]. This task is evaluated in both zero-shot and by fine-tuning the decoder, reporting Pearson correlation between predicted and true gene expression values. We explain these tasks in further detals in Appendix C.2.
Baselines. We compare HEIST with graph-based spatial models STAGATE [9] and GraphST [26], which focus on local spatial relationships but lack generalization across datasets, and with scFoundation [18], a large-scale single-cell foundation model that ignores spatial context. We also include comparisons with recent spatial foundation models CellPLM [39] and scGPT-spatial [38], which incorporate spatial information but do not explicitly model hierarchical gene-cell interactions. These baselines provide a comprehensive evaluation across specialized spatial models, non-spatial foundation models, and spatial foundation models, demonstrating HEIST’s ability to generalize across biological scales and outperform both specialized and general-purpose models.
Figure 3: HEIST accounts for tissue microenvironments. A. Comparison of cell embeddings for the same tissue slice from SEA colored by cell types. HEIST demonstrates superior separation of distinct cell types compared to other methods. B. Spectral clustering of HEIST embeddings reveals sub-clustering within cell types (refer to Appendix  C.2 for more details). C. L4-IT neurons investigation shows that clusters accounts for spatial organization and neighborhood differences, demonstrating that HEIST embeddings effectively captures spatial microenvironments through cross-message passing. Numerical results are available in Table A3 in Appendix Section D.
4.1 Results
Spatially-aware cell type discovery. Figure 3 shows that HEIST is able to do separate spatially-aware subclusters based on cell-embeddings, while existing foundation models fail to do so. Our model separates known cell types and further reveals microenvironment-driven subpopulations, enabling the discovery of new, spatially informed cell types. Meanwhile, CellPLM an scGPT-spatial fail to address the microenvironment separation as shown in Figure A2 in Appenndix.
Tissue-classification. Table 1 reports the tissue-level classification performance of various models across multiple datasets and tasks, measured by AUC-ROC. HEIST consistently outperforms existing models, achieving the highest scores in six out of seven evaluation scenarios. Notably, on the UPMC tasks, HEIST surpasses the foundation models scGPT-spatial and CellPLM by 25.4% and 30%, respectively. Since CODEX is not a technology that scGPT-spatial and CellPLM were trained on, they support only 6 out of 40 genes in the UPMC datasets. As a result, these models ignore most of the available gene information, leading to significant performance degradation. In contrast, HEIST leverages GRNs and models gene dynamics directly, rather than relying on a fixed set of genes. This enables HEIST to achieve substantially better performance across diverse technologies and datasets.
Organ Colon Neck Neck Skin Placenta
Dataset Charville UPMC DFCI Melanoma
Model Task Outcome Recurrence Outcome Recurrence Outcome Response Condition
STAGATE 0.657 $\pm$ 0.032 0.783 $\pm$ 0.050 0.602 $\pm$ 0.054 0.659 $\pm$ 0.013 0.633 $\pm$ 0.210 0.533 $\pm$ 0.267 0.578 $\pm$ 0.12
GraphST 0.659 $\pm$ 0.059 0.828 $\pm$ 0.088 0.645 $\pm$ 0.026 0.582 $\pm$ 0.061 0.683 $\pm$ 0.131 0.567 $\pm$ 0.170 0.644 $\pm$ 0.15
Space-gm* 0.793 0.696 0.863 0.883 0.873 - -
[…]
CellPLM 0.744 $\pm$ 0.006 0.801 $\pm$ 0.032 0.681 $\pm$ 0.134 0.667 $\pm$ 0.045 0.750 $\pm$ 0.083 0.580 $\pm$ 0.133 0.682 $\pm$ 0.15
scGPT-spatial 0.834 $\pm$ 0.081 0.806 $\pm$ 0.019 0.717 $\pm$ 0.117 0.676 $\pm$ 0.097 0.875 $\pm$ 0.041 0.600 $\pm$ 0.000 0.602 $\pm$ 0.08
HEIST 0.861 $\pm$ 0.086 0.887 $\pm$ 0.041 0.835 $\pm$ 0.001 0.929 $\pm$ 0.030 0.937 $\pm$ 0.062 0.866 $\pm$ 0.066 0.769 $\pm$ 0.06
HEIST Imp. % 3.2 10.0 -3.3 5.2 7.1 44.3 12.7
Table 1: Performance on clinical outcome prediction. We classify cancer outcome, cancer remission, treatment response and placental conditions. *Results are taken from [41]
Cell-type annotation. Table 2 shows the performance on the cell type annotation task across multiple datasets. HEIST achieves the highest performance in four out of five datasets, with substantial gains in UPMC and DFCI neck datasets (311.3% and 81.9% improvements, respectively). Notably, scGPT-spatial was pretrained on the MERFISH Lung Cancer dataset, explaining its strong performance there, while HEIST remains competitive despite not being trained on that data. These results again highlight the effectiveness and generalizability of HEIST.
Organ Brain Colon Neck Neck Lung
Model SEA-AD Charville UPMC DFCI
STAGATE 0.3304 $\pm$ 0.0625 0.2759 $\pm$ 0.0490 0.0687 $\pm$ 0.0136 0.0685 $\pm$ 0.0213 0.2187 $\pm$ 0.0570
GraphST 0.2296 $\pm$ 0.1772 0.3675 $\pm$ 0.0873 0.0617 $\pm$ 0.0199 0.0577 $\pm$ 0.0261 0.4081 $\pm$ 0.0658
[…]
CellPLM 0.6701 $\pm$ 0.0827 0.4760 $\pm$ 0.0669 0.0563 $\pm$ 0.0212 0.0565 $\pm$ 0.0179 0.5044 $\pm$ 0.1607
scGPT-spatial 0.5907 $\pm$ 0.0029 0.3494 $\pm$ 0.0624 0.0464 $\pm$ 0.0162 0.0618 $\pm$ 0.0163 0.5671 $\pm$ 0.1685
HEIST 0.9953 $\pm$ 0.0158 0.5340 $\pm$ 0.1293 0.2826 $\pm$ 0.0758 0.1124 $\pm$ 0.0521 0.5126 $\pm$ 0.1170
HEIST Imp. % 48.5 12.2 311.3 81.9 -9.6
Table 2: Performance on cell type annotation. Annotations are provided as a feature in each dataset.
Model Placenta Skin
ScFoundation (Fine-tuned) 0.721 $\pm$ 0.004 0.621 $\pm$ 0.003
CellPLM (Fine-tuned) 0.801 $\pm$ 0.011 0.723 $\pm$ 0.007
scGPT-spatial (Fine-tuned) 0.718 $\pm$ 0.002 0.740 $\pm$ 0.002
HEIST (Zero-Shot) 0.574 $\pm$ 0.000 0.350 $\pm$ 0.000
HEIST (Fine-tuned) 0.821 $\pm$ 0.041 0.807 $\pm$ 0.020
HEIST Imp. % 2.49 9.05
Table 3: Performance on gene imputation task
Gene imputation. Table 3 reports the performance on the gene imputation task for the placenta and skin datasets. HEIST achieves the best performance after fine-tuning, surpassing all baseline models by 2.5% on the placenta dataset and 9.1% on the skin dataset. Although HEIST’s zero-shot performance is limited by the dataset-specific nature of gene expression patterns, fine-tuning allows the model to effectively adapt by leveraging its hierarchical structure and GRN-based representations, resulting in improved performance.
Ablation of components. Table 4 shows that removing any key component of HEIST leads to performance degradation across tasks, confirming their importance. Hierarchical modeling and spatial information are most critical, with their removal causing the largest drops. Cross-level message passing and contrastive learning significantly improve cell classification, while the masked autoencoder (MAE) is crucial for gene imputation and classification. Orthogonal regularization mainly benefits gene imputation but has less impact on other tasks.
Model Charville-Outcome Skin-Imputation SEA-Cell classification
HEIST 0.861 $\pm$ 0.086 0.807 $\pm$ 0.020 0.995 $\pm$ 0.015
No space (No Hierarchy) 0.596 $\pm$ 0.028 0.345 $\pm$ 0.010 0.179 $\pm$ 0.038
No gene (No Hierarchy) 0.764 $\pm$ 0.235 0.173 $\pm$ 0.014 0.194 $\pm$ 0.040
No cross message passing 0.625 $\pm$ 0.125 0.531 $\pm$ 0.005 0.955 $\pm$ 0.041
No positional encodings 0.523 $\pm$ 0.010 0.458 $\pm$ 0.003 0.220 $\pm$ 0.034
No contrastive 0.623 $\pm$ 0.002 0.536 $\pm$ 0.015 0.966 $\pm$ 0.037
No MAE 0.658 $\pm$ 0.076 0.495 $\pm$ 0.006 0.162 $\pm$ 0.038
No orthogonal regularization 0.594 $\pm$ 0.031 0.646 $\pm$ 0.013 0.992 $\pm$ 0.020
Table 4: Ablation study showing that hierarchical modeling, cross-level message passing, and the training objectives are critical components for strong performance.
5 Conclusion
We present HEIST, the first hierarchical foundation model for spatial transcriptomics that jointly models gene regulatory networks and spatial cell-cell interactions within a unified framework. By integrating biologically motivated hierarchical representation learning with a novel cross-level message passing mechanism, HEIST captures complex dependencies between genes, cells, and tissue-level organization. Pretrained on over 22.3 million cells spanning 15 organs, HEIST achieves state-of-the-art performance across multiple downstream tasks and technologies, while offering significant computational efficiency. Beyond predictive accuracy, HEIST enables the discovery of spatially-informed cellular subpopulations, providing deeper insights into tissue microenvironments. Our results highlight the importance of modeling both molecular and spatial hierarchies to advance the development of general-purpose, transferable models for biological systems. While HEIST represents a significant step toward general-purpose, transferable models for biological systems, it does not yet capture temporal tissue dynamics, pointing to promising future directions. Nonetheless, by enabling new insights into tissue organization and supporting open science, HEIST has the potential to accelerate discoveries in areas such as cancer research and precision medicine. A detailed discussion of limitations and societal impact is provided in Appendix Sections E and F, respectively.
Acknowledgments
S.K. is funded in part by the NIH (NIGMSR01GM135929, R01GM130847), NSF CAREER award IIS-2047856, NSF IIS-2403317, NSF DMS-2327211. S.K is also funded by the Sloan Fellowship FG-2021-15883, the Novo Nordisk grant GR112933. S.K. and R.Y. acknowledge funding from NSF CISE-2403317. R.Y. has also recieved Amazon Research Award 2024.
Appendix A Data Preprocessing and Graph creation
HEIST requires construction of two graph structures: (1) a spatial cell-cell graph capturing local tissue organization, and (2) cell-type-specific graph representing the gene regulatory networks (GRNs), which captures gene-gene dependencies conditioned on cell types. We present the preprocessing pipeline in Figure A1, and we describe the data preprocessing steps and the detailed procedures for constructing these graphs below.
Data Preprocessing. We begin by importing raw spatial transcriptomics data, followed by standard preprocessing steps, including outlier removal, gene expression normalization, and gene filtering to retain highly variable genes. To mitigate technical noise and dropout effects common in single-cell measurements, we apply MAGIC [8] to denoise gene expression data before all downstream computations.
Cell-Cell Graph Construction. To capture spatial relationships between cells, we construct a cell-cell graph based on physical proximity. Spatial locations are used to compute Voronoi polygons, which define local neighborhoods. Edges in the graph are assigned based on adjacency in the Voronoi diagram, ensuring that the graph structure accurately reflects tissue architecture and local microenvironments.
Cell-Type-Specific GRN Construction. For each cell type, we build a gene regulatory network (GRN) that captures functional dependencies between genes. If cell type annotations are provided, we use them directly; otherwise, we infer cell types via Leiden clustering on the denoised gene expression data. After cell types are assigned, we compute pairwise mutual information (MI) between genes within each cell type using the denoised expression values. Gene pairs with MI greater than a threshold $\tau$ are connected by an edge in the GRN, resulting in a sparse, cell-type-specific regulatory graph. This process captures gene-gene co-expression patterns and regulatory dependencies unique to each cellular context.
Final Output. The final preprocessing pipeline produces:
  • A spatial ${\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}})$ cell-cell graph encoding local tissue structure.
  • Cell-type-specific GRNs $\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}$ capturing regulatory function within each cell type.
These graph structures provide the foundation for HEIST’s hierarchical learning framework, enabling integration of spatial organization and gene regulatory patterns.
Figure A1: HEIST preprocessing pipeline
Appendix B Implementation details
In this section, we explain the implementation details including the positional encoding initialization, graph transformer architecture, and hyperparameters used in pre-training phase.
B.1 Positional encodings
HEIST incorporates positional encodings (PE) at both the cell and gene levels to inject spatial and regulatory structure into the learned representations. As opposed to traditional graph PEs such has Laplacian PE [10], random walk PE [10], node-centrality based PE [44], we make use of sinusoidal PEs. This design choice is motivated by the fact that sinusoidal PE yields expressive representations and are computationally in-expensive as opposed to traditioanl graph PEs, leading to efficient and expressive representations [40]. Below, we describe how these encodings are constructed and why they are suitable for spatial transcriptomics.
Each cell is associated with a spatial coordinate $(x,y)\in\mathbb{R}^{2}$, corresponding to its location in the tissue. To encode spatial information, we apply a two-dimensional extension of the sinusoidal positional encoding introduced in transformers [35]:
$\text{PE}_{c,2i}$ $=\sin\left(\frac{x}{10000^{4i/d}}\right),\quad\text{PE}_{c,2i+1}=\cos\left(\frac{x}{10000^{4i/d}}\right),$
$\text{PE}_{c,2j+d/2}$ $=\sin\left(\frac{y}{10000^{4j/d}}\right),\quad\text{PE}_{c,2j+1+d/2}=\cos\left(\frac{y}{10000^{4j/d}}\right),$
where $d$ is the dimensionality of the encoding. To calculate the gene-PE, for each cell, genes are first sorted in descending order of expression. Since most of the spatial transcriptomics are noisy, ranking is done after denoising using MAGIC. The rank of each gene $g$ reflects its relative expression within the cell:
$\text{Rank}_{k}(g)=\text{position of }g\text{ in sorted list by expression of cell $k$}.$
We then apply sinusoidal encoding based on the rank values as follows:
$\text{PE}^{k}_{g,2i}=\sin\left(\frac{\text{Rank}_{k}(g)}{10000^{2i/d}}\right),\quad\text{PE}^{k}_{g,2i+1}=\cos\left(\frac{\text{Rank}_{k}(g)}{10000^{2i/d}}\right).$
This treats genes within a cell as a soft sequence, allowing the model to distinguish highly expressed genes from lowly expressed ones based on their position in the transcriptional program. Sinusoidal encodings on rank preserve relative ordering and allow the model to capture patterns in gene expression. Since the same gene can have different ranks across cells, the resulting embedding is context-dependent, enabling the model to learn cell-specific regulatory roles of genes. This is important for spatial transcriptomics, where the function and relevance of a gene may vary depending on the cell type and microenvironment. Together, the PE helps HEIST incorporate spatial geometry and transcriptional ordering in a biologically meaningful and computation-efficient way.
B.2 Intra-level message passing
We incorporate biological inductive biases by applying global (all-to-all) attention on the cell graph to capture long-range spatial interactions and local attention that depends on the graph structure of gene graphs to model fine-grained regulatory patterns [29]. Capturing global dependencies is crucial for detecting large-scale tissue structures, such as immune infiltration zones and tumor-stroma boundaries [47], while modeling local regulatory networks enables identification of tightly coordinated gene modules within specific niches [17]. HEIST performs intra-level message passing independently at the cell and gene levels to capture both local and global dependencies within each modality.
At the cell level, we compute:
$\tilde{{\mathbf{H}}}_{c}^{(l)}=\text{GraphTransformerLayer}({\mathbf{H}}_{c}^{(l-1)})=\text{TransformerLayer}({\mathbf{H}}_{c}^{(l-1)})+\text{GIN}({\mathbf{A}}_{c},{\mathbf{H}}_{c}^{(l-1)}),$
where ${\mathbf{H}}_{c}^{(l)}$ denotes the cell embeddings at layer $l$, and ${\mathbf{A}}_{c}$ is the adjacency matrix of the spatial cell graph. The transformer layer enables long-range interactions across spatial regions, while the GIN layer aggregates local neighborhood information. This hybrid approach captures both global spatial context (e.g., tissue-level structure) and localized microenvironments (e.g., niches or boundaries).
At the gene level, we apply sparse attention based on the connectivity of the genes as:
$\alpha_{ij}^{k^{(l)}}=\frac{(\mathbf{W}_{q}{{\mathbf{h}}}_{g,i}^{k^{(l-1)}})^{\top}(\mathbf{W}_{k}{{\mathbf{h}}}_{g,j}^{k^{(l-1)}})}{\sqrt{d}},\quad\text{if}(i,j)\in\mathcal{E}_{k},$
$\tilde{{{\mathbf{h}}}}_{g,i}^{k^{(l)}}=\sum_{j\in\mathcal{N}_{g}^{k}(i)}\text{softmax}_{j}(\alpha_{ij}^{k^{(l)}})\mathbf{W}_{v}{{\mathbf{h}}}_{g,j}^{k^{(l-1)}},$
where $\mathbf{h}_{g,i}^{k^{(l)}}$ is the embedding of gene $i$ in cell $k$ at layer $l$, and $\mathcal{N}_{g}^{k}(i)$ denotes the GRN neighbors of gene $i$ in that cell. This formulation ensures attention is computed only over biologically relevant interactions, preserving regulatory sparsity while allowing expressive, cell-specific dependency modeling.
B.3 Hyperparameters
In Table A1, we describe the hyperparameters for the pre-training model.
Hyperparameter Default Value
Positional encoding dimension 128
Hidden dimension 128
Output dimension 128
Number of HEISTLayer layers 10
Number of transformer heads 8
Batch size 256
Learning rate 0.001
Weight decay 0.003
Number of pre-training epochs 20
Activation function GeLu [19]
Optimizer AdamW [27]
Table A1: Hyperparameters and their default values used in the experiments.
Appendix C Datasets
In this section, we first briefly talk about pre-training datasets, followed by downstream tasks explained in details and downstream datasets.
C.1 Pretraining dataset
The pre-training dataset of HEIST comprised of 22.3 million cells from 124 tissues slices and 15 organs, including data collected through MERFISH and Xenium. The data consists of 13.3M cells from 10xGenomics datasets (https://www.10xgenomics.com/), 8.7M from Vizgen (https://vizgen.com/) and 360k cells from Seattle Alzheimer’s Brain Atlas (https://portal.brain-map.org/explore/seattle-alzheimers-disease).
In contrast to previous models that incorporate lower-resolution data from Visium arrays [7], HEIST focuses exclusively on single-cell resolution datasets. This design choice reflects the fact that Visium data lacks true single-cell resolution, instead capturing transcriptomic signals aggregated over spatial spots that often contain heterogeneous mixtures of multiple cell types. Such coarse-resolution data can introduce confounding signals and limit the model’s ability to accurately learn cell-level spatial dependencies and gene regulatory programs. By focusing on single-cell spatial transcriptomics, HEIST is better positioned to capture fine-grained spatial organization, cell-cell interactions, and context-dependent gene regulation critical for modeling tissue microenvironments and cellular heterogeneity.
C.2 Downstream Tasks
In this section, we explain the tasks in more details
Cell type annotation. This task involves assigning cells to known cell types based on their gene expression and spatial context. We evaluate this by extracting representations from a frozen HEIST encoder and training only an MLP head for classification. Accurate cell type annotation is essential for characterizing cellular composition across tissues and studying how cell populations contribute to tissue function and disease.
Cell clustering. Cell clustering is performed in an unsupervised manner using frozen HEIST embeddings to discover both known and novel cellular subpopulations. Unlike prior models, HEIST’s spatially informed representations not only separate cells into canonical cell types but also induce subclusters that reflect the influence of local microenvironments. This enables the discovery of previously unrecognized, spatially contextualized cell types based on their tissue niche and surrounding cellular interactions. As shown in Figure 3, HEIST effectively identifies such microenvironment-specific subpopulations, which existing foundation models fail to capture. Figure A2 does the same analysis on scGPT-spatial and CellPLM showing that these methods fail to address the microenvironment separation.
Figure A2: Comparisons between embeddings focusing on L4-IT neurons. The figure shows the cell type embeddings, followed by spectral clustering and a plot of all clusters that contain the specific cell type. A. scGPT-spatial results. B. CellPLM results C. HEIST results (same as Figure 3) D. Ground Truth cell types colored in space.
Tissue-level classification. In this task, the goal is to predict tissue-level outcomes, including response to immunotherapy, treatment outcomes, remission status, and placenta condition. Accurate classification requires capturing both local cellular microenvironments and broader tissue organization patterns. We first evaluate HEIST on datasets from Charville (Colon), UPMC (Neck), and DFCI (Neck), collected using CO-Detection by Indexing (CODEX) technology. We also predict immunotherapy response on a skin cancer dataset [30] collected through multiplex ion beam imaging (MIBI). Finally, we classify placenta condition into normal placenta, placenta accreta spectrum (PAS), and placental insufficiency using data collected with Xenium. Spatial modeling is crucial for placenta classification, as conditions like PAS involve disrupted tissue architecture and abnormal cell invasion, which gene expression alone cannot fully capture. For this task, we extract cell representations from a frozen HEIST encoder and train an MLP head. We report AUC-ROC as the evaluation metric.
Gene imputation. In spatial transcriptomics, gene expression profiles are often incomplete or noisy due to technical limitations and measurement dropouts. Gene imputation aims to predict missing gene expression values, recovering biologically plausible expression patterns. HEIST performs this task by leveraging the learned gene embeddings, which capture both regulatory relationships from GRNs and spatial dependencies from their parent cells. Through cross-level message passing, the model can incorporate spatial cues to refine gene predictions, enabling more accurate reconstruction of missing data. We select the genes to be masked using stratified sampling according to gene sparsity [1] and mask these genes during the fine-tuning phase. This task can be performed in a zero-shot fashion or by fine-tuning the HEIST decoder on the downstream task. We report Pearson correlation as the evaluation metric for this task.
C.3 Downstream datasets
We have curated different datasets for each downstream prediction task. To show our models generalize well we specifically used 4 different technologies across the different tasks. The dataset description can be found on Table A2
Dataset Organ # Slices # Cells Technology Tissue Cell-type Gene
Classification Annotation Clustering imputation
Lung Cancer Lung 1 100,000 MERFISH $\times$ $\times$
DFCI Head and neck 58 125,512 CODEX $\times$
UPMC Head and neck 308 2,164,932 CODEX $\times$
Charville Colon 292 632,180 CODEX $\times$
Melanoma Skin 54 540,000 MIBI $\times$ $\times$
Placenta Placenta 212 1,000,000 Xenium $\times$ $\times$
Table A2: Summary of spatial-omics datasets used in this study. For each dataset, we indicate the organ of origin, number of tissue slices analyzed, total cell count, imaging technology used, and the tasks performed.
Appendix D Extra results
Cell clustering. Table A3 reports normalized mutual information (NMI) scores for the cell clustering task across multiple organs and datasets. HEIST achieves the highest or comparable NMI scores in most settings, demonstrating its ability to capture meaningful cellular subpopulations informed by both spatial and regulatory contexts. Notably, HEIST shows substantial improvement on datasets with complex tissue structures, such as UPMC and DFCI, where modeling spatial microenvironments is critical. While some models achieve competitive performance in simpler settings, they fail to generalize as effectively across technologies due to missing genes in their gene set.
Organ Brain Colon Neck Neck Lung
Model SEA-AD Charville UPMC DFCI MERFISH Lung cancer
STAGATE 0.294 $\pm$ 0.05 0.237 $\pm$ 0.04 0.022 $\pm$ 0.01 0.048 $\pm$ 0.02 0.171 $\pm$ 0.06
GraphST 0.444 $\pm$ 0.02 0.224 $\pm$ 0.05 0.020 $\pm$ 0.01 0.044 $\pm$ 0.02 0.297 $\pm$ 0.05
[…]
CellPLM 0.651 $\pm$ 0.00 0.24 $\pm$ 0.00 0.015 $\pm$ 0.00 0.075 $\pm$ 0.00 0.286 $\pm$ 0.00
scGPT-spatial 0.674 $\pm$ 0.04 0.253 $\pm$ 0.07 0.017 $\pm$ 0.01 0.038 $\pm$ 0.01 0.307 $\pm$ 0.08
HEIST 0.691 $\pm$ 0.04 0.297 $\pm$ 0.03 0.043 $\pm$ 0.01 0.14 $\pm$ 0.04 0.274 $\pm$ 0.04
Table A3: Comparison of unsupervised cell clustering performance using Normalized Mutual Information (NMI).
Runtime comparison. Table A4 compares the runtime efficiency of HEIST with existing foundation models for generating embeddings from a tissue sample with 19,826 cells. HEIST achieves substantial speed improvements, being 8× faster than scGPT and 48× faster than scFoundation, while providing both cell and gene embeddings. Although CellPLM is faster, it does not compute gene representations, making it less suitable for tasks requiring joint cell and gene analysis.
Model Runtime
scFoundation 290.48
scGPT 49.37
HEIST 6.69
CellPLM* 1.97
Table A4: Runtime (in seconds) comparison to extract embeddings from a tissue with 19826 cells. * CellPLM does not compute gene representations.
PE ablation. Table A5 presents an ablation study comparing sinusoidal PE against traditional graph-based PEs—random walk and Laplacian—on clinical outcome prediction task. Sinusoidal PE consistently yields higher accuracy and lower variance, particularly excelling in recurrence prediction, highlighting its effectiveness in capturing spatial patterns relevant to clinical signals.
Organ Colon Neck Neck
Dataset Charville UPMC DFCI
Model Task Outcome Recurrence Outcome Recurrence Outcome
Random walk PE 0.770 $\pm$ 0.034 0.695 $\pm$ 0.042 0.776 $\pm$ 0.029 0.752 $\pm$ 0.038 0.916 $\pm$ 0.026
Laplacian PE 0.750 $\pm$ 0.031 0.800 $\pm$ 0.025 0.670 $\pm$ 0.045 0.810 $\pm$ 0.021 0.852 $\pm$ 0.040
Sinusoidal PE 0.861 $\pm$ 0.086 0.887 $\pm$ 0.041 0.835 $\pm$ 0.001 0.929 $\pm$ 0.030 0.937 $\pm$ 0.062
Table A5: Ablation comparing sinusoidal PE based on spatial coordinates with traditional graph-based PEs for clinical outcome prediction.
Appendix E Limitations
While HEIST advances the state of foundation models for spatial transcriptomics (ST), it has several limitations. First, the current GRN construction relies on co-expression relationships using mutual information, which may not fully capture causal gene regulatory mechanisms or directional influences. This can be integrated by more sophisticated MI meassures like DREMI [23]. Furthermore, integrating more sophisticated GRN inference techniques could improve the biological interpretability and efficiency of gene embeddings. A potential direction for future work can be incorporating temporal dynamics by applying techniques such as Granger causality [34] over inferred pseudotime trajectories, allowing the model to capture not only static spatial and regulatory dependencies but also directional gene regulatory influence and developmental progression, which are currently not modeled in HEIST. Second, the model assumes static spatial snapshots of tissues and does not account for temporal dynamics or developmental trajectories, which are critical in understanding certain biological processes. Extending HEIST to model spatio-temporal transcriptomics data is an important direction for future work. Finally, while HEIST improves computational efficiency over prior foundation models, it still requires substantial computational resources for large-scale pretraining. Despite these limitations, HEIST provides a flexible and scalable foundation for modeling complex spatial and molecular interactions, and future work can address these challenges to further improve its generalization and interpretability.
Appendix F Broader Impacts
HEIST advances the state of foundation models for spatial transcriptomics, providing powerful tools for understanding complex gene regulation and tissue organization. This has the potential to accelerate scientific discovery in areas such as cancer research, immunotherapy, and developmental biology, and to support the design of more effective precision medicine strategies. By enabling better modeling of microenvironments and gene regulatory dynamics, HEIST contributes to uncovering novel biomarkers and therapeutic targets. The model is trained entirely on publicly available, anonymized biological datasets, with no use of personal or identifiable human data, minimizing concerns related to privacy or misuse. While HEIST produces biologically meaningful embeddings, we emphasize that it is a tool to support and augment expert biological analysis, rather than replace it. Experimental validation remains essential for translating model insights into clinical or therapeutic applications. By releasing our code and models publicly, we aim to foster open science, reproducibility, and collaborative advancements in computational biology. We are optimistic that HEIST will serve as a valuable resource for the scientific community, contributing positively to both research and education in systems biology and bioinformatics.